OpenStudy (anonymous):

What is the sum of the roots of the polynomial shown below? ƒ(x)=x^3-8x^2-23x+30

OpenStudy (anonymous):

look at the coefficient of the power of x to the right of the leading power of x. its the negative of that. (lol that sounds confusing)

OpenStudy (anonymous):

The second coefficient in that polynomial is -8 so the sum of the roots is 8

myininaya (myininaya):

f(1)=0 so x=1 is a zero => x-1 is a factor 1 | 1 -8 -23 30 | 1 -7 -30 --------------------- 1 -7 -30 | 0 x^2-7x-30 is another factor x^2-7x-30=(x-12)(x+5) x=12, x=-5 our roots too s0 1+12+(-5)=13-5=8

myininaya (myininaya):

but that looks faster what joe has lol

myininaya (myininaya):

i never seen this way joe

OpenStudy (anonymous):

if you ever have a polynomial like: \[x^n+a_{n-1}x^{n-1}+a_{n-2}x^{n-2}+\cdots +a_1x+a_0 = 0\] the sum of the roots is: \[-a_{n-1}\] and the product of the roots is: \[(-1)^na_0\]

myininaya (myininaya):

wow i like this can we prove it?

OpenStudy (anonymous):

Would you like to do the proof form scratch? I think i have a proof lying around on my computer cause I had to do it for a test once.

myininaya (myininaya):

i'm going to watch a movie and i want to try it on my own so don't post anything k?

OpenStudy (anonymous):

sure thing :)

myininaya (myininaya):

see ya later :)